Find the LCM of x^2+6x+9 and x^2+x-6
Answer by nerdybill(7384)

Find the LCM of x^2+6x+9 and x^2+x-6
.
factoring:
x^2+6x+9 = (x+3)(x+3)
.
x^2+x-6 = (x+3)(x-2)
.
LCM:
(x+3)(x+3)(x-2)